# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ 3.1 Data and assumptions
As per Section 2, let $T=t$ denote a time-invariant binary treatment, with $t\in\{0,1\}$, such that $T=1$ represents the active intervention and $T=0$ the control. Let $S=s$ denote the data source, with $s\in\{0,1\}$, such that $S=1$ represents the SAT and $S=0$ the external data source. In addition, let $\mathbf{X}$ denote vector-valued pre-treatment baseline covariates, e.g., clinical or demographic characteristics, measured across the SAT and the external data source. Let $Y$ denote the clinical outcome of interest. We assume that only distributional differences in $\mathbf{X}$ are preventing exchangeability between the SAT and external subjects, and that covariates and outcomes are defined and measured similarly across data sources.
The observed IPD consist of $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$, $i=1,\dots,n=n_{1}+n_{0}$, realizations of $(S,\mathbf{X},T,Y)$ denoting the data source, baseline covariates, treatment assignment and observed outcome for subject $i$. Here, the SAT and external data source have been stacked, with $n_{1}$ and $n_{0}$ as the sample sizes of the SAT and the external data source, respectively. It is assumed that all individuals in the SAT are under $T=1$ and all individuals in the external data source are under $T=0$, such that the control group is fully external. To be clear, we have $S_{i}=T_{i}$ for all $i=1,\dots,n$; and also have that $S_{i}=1$ and $T_{i}=1$ for all $i=1,\dots,n_{1}$, and $S_{i}=0$ and $T_{i}=0$ for all $i=n_{1}+1,\dots,n$. We shall assume that there is no missingness or measurement error.
The observed outcome for subject $i$ is $Y_{i}=Y_{i}^{1}T_{i}+Y_{i}^{0}(1-T_{i})$, where $Y_{i}^{t}$ is the potential outcome had subject $i$ been assigned treatment $t\in\{0,1\}$, with $Y_{i}=Y_{i}^{1}$ if $i=1,\dots,n_{1}$ and $Y_{i}=Y_{i}^{0}$ if $i=n_{1}+1,\dots,n$. Namely, the observed outcome for an individual in the SAT equals their potential outcome under the active intervention, and the observed outcome for an individual in the external data source equals their potential outcome under the control. Implicit in the notation is the stable unit treatment value assumption (SUTVA): that there is no interference between subjects and there is treatment version irrelevance, i.e., one well-defined version of the active intervention and the control across all subjects and data sources.54 Also implicit is that there is no direct effect of trial participation.54 Namely, that trial participation – in the SAT or a historical trial, for that matter – does not affect the outcome except through treatment assignment itself, i.e., there are no Hawthorne effects.55, 56
@@
[…]
To enforce positivity, analysts may subset the SAT based on the selection criteria of the external control.25, 28 However, this further reduces the sample size of the SAT. Positivity is typically assessed by comparing the empirical distributions of the covariates in the SAT and the external control.60 While outcome modeling-based approaches such as G-computation can overcome failures of positivity, they do so by potentially problematic and difficult-to-diagnose model-based extrapolation. Even minor model misspecification over the observed covariate space in the SAT may lead to poor extrapolation in unobserved regions.48
Analogously, targeting the ATT would require constructing estimators for $\mu^{1}_{1}=\textrm{E}(Y^{1}\mid S=1)$ and $\mu^{0}_{1}=\textrm{E}(Y^{0}\mid S=1)$. Here, the challenge is the estimation of $\mu^{0}_{1}$ because outcomes under the control have not been generated in the SAT. The conditional ignorability assumption would formally be $Y_{i}^{0}\perp\!\!\!\!\perp S_{i}\mid\mathbf{x}_{i}$ for all $i=1,\dots,n$, and would rest on the SAT and the external control measuring all variables that are prognostic of outcome under the control. The positivity assumption would be $10<\textrm{Pr}(S=0\mid\mathbf{X}=\mathbf{x})<$ 1 for all $\mathbf{x}$ with positive density in the SAT, $f(\mathbf{x}\mid S=1)>0$, such that the support of the baseline covariates in the SAT is contained within that of the external control and it is possible to have external control subjects in all regions of the SAT covariate distribution. In this setting, analysts may apply the SAT selection criteria to the external control to guarantee that there is sufficient overlap.
@@ 3.2 Inverse odds weighting
We first present a covariate adjustment method that models the data source assignment mechanism, conditional on baseline covariates, to estimate weights.33, 34, 40, 61, 62 Where the target estimand is the ATC, SAT subjects are weighted by their inverse conditional odds of SAT participation – their conditional odds of external control participation – to transport the SAT outcomes to the external control (sub) population. Such “inverse odds” weights (IOW) are defined as:
@@
where $\operatorname{logit}(e_{i})=\ln\left((e_{i})/(1-e_{i})\right)$, $\alpha_{0}=\ln\left(\textrm{Pr}(S_{i}=1|\mathbf{c}(\mathbf{X}_{i})=0)/(\textrm{Pr}(S_{i}=0|\mathbf{c}(\mathbf{X}_{i})=0)\right)$ is an intercept term, $\mathbf{\alpha}$ is a vector of regression parameters, and $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i}),\dots,c_{k}(\mathbf{X}_{i})]^{\top}$ is a vector of covariate “balance functions” for subject $i=1,\dots n$. This is the set of functions containing the distributional features to be balanced between the SAT and the external control,40, 63 potentially including sensible transformations of the covariates, e.g., polynomials and interaction terms.
The logistic regression is typically fitted to the concatenated IPD using maximum-likelihood estimation, with the regression coefficient point estimates denoted by $\hat{\alpha}_{0}$ and $\hat{\mathbf{\alpha}}$, and model-based propensity scores for subject $i=1,\dots,n_{1}$ predicted by $\hat{e}_{i}=\operatorname{logit}^{-1}\left(\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\alpha}}\right)=\operatorname{expit}\left(\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\alpha}}\right)$, where $\operatorname{expit}(\cdot)=\exp(\cdot)/\left(1+\exp(\cdot)\right)$. Weight estimates $\hat{w}_{i}$ for $i=1,\dots,n_{1}$ are derived by plugging the corresponding propensity score predictions into Equation 1. With correct specification of the model in Equation 2, such that the log-odds of SAT participation are linear across the balance functions of the covariates, $\hat{e}_{i}$ and $\hat{w}_{i}$ consistently estimate the true conditional probability and inverse odds of SAT participation, respectively.
The ATC is estimated by contrasting the weighted average of observed outcomes under the active intervention with the unweighted average of observed outcomes for the external control. As per Section 2, mean absolute outcomes are converted to the additive scale imposed by link function $g(\cdot)$ prior to taking the difference between treatments on such scale:
@@
The mean absolute outcome estimate for the active intervention can be bounded within its feasible range, e.g., between 0 and 1 for probabilities, by normalizing or “stabilizing” the weights so that they sum to one.57, 61 This results in the alternative ATC estimator:
@@
which should provide improved finite sample properties and more stable and precise estimation.64, 65
In expectation, if the model in Equation 2 is correctly specified, the estimated weights, ($\hat{w}_{i}$, for $i$ in $n_{1}$) will balance the covariate distribution of the SAT with respect to that of the external control, enabling consistent estimation of mean absolute outcome $\mu^{1}_{0}$ and the ATC. To see that the IOW estimators are consistent, consider a simple scenario with a binary outcome, $Y$, and a single discrete covariate, $X$, such that:
$\mu_{0}^{1}=\textrm{E}(Y|T=1,S=0)$ $=1\times\textrm{Pr}(Y=1|T=1,S=0)+{0\times\textrm{Pr}(Y=0|T=1,S=0)}$
$=\textrm{Pr}(Y=1|T=1,S=0).$
Basic probability rules imply that the marginal risk is the weighted average of the stratum-specific risks:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,S=0,X=x)\textrm{Pr}(X=x|S=0,T=1).$
Then, due to the assumption of conditional data source ignorability, we have:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=0)$
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=0)\frac{\textrm{Pr}(X=x|S=1)}{\textrm{Pr}(X=x|S=1)}$
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(X=x|S=0)}{\textrm{Pr}(X=x|S=1)},$
and from Bayes’ rule, we have:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}\frac{\textrm{Pr}(S=1)}{\textrm{Pr}(S=0)}$
$=\underbrace{\frac{\textrm{Pr}(S=1)}{\textrm{Pr}(S=0)}}_{A}\underbrace{\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}}_{B}.$ (5)
Since $\sum_{i=1}^{n}{\frac{S_{i}}{n}}=n_{1}/n\rightarrow\textrm{Pr}(S=1)$ and $\sum_{i=1}^{n}{(1-S_{i})}/n=n_{0}/n\rightarrow\textrm{Pr}(S=0)$, we can consistently estimate $A\approx{n_{1}}/{n_{0}}$. We can also consistently estimate $B$ from the sample since the covariate distribution for the SAT, $(X|S=1)$, is observed (and since, within the sample, $T=1\iff S=1$):
$B$ $\approx\sum_{i=1}^{n_{1}}\textrm{Pr}(Y=1|S=1,X=x_{i})\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}\approx\sum_{i=1}^{n_{1}}\frac{Y_{i}}{n_{1}}\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}.$
Finally, if the propensity score model is correctly specified, the inverse odds weights consistently estimate the true inverse odds, such that, for $i$ in $n_{1}$:
$\widehat{w}_{i}\rightarrow$ $\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}.$ (6)
Therefore, the IOW weighted estimator is consistent:
$\hat{\mu}_{0}^{1}$ $=\underbrace{\frac{n_{1}}{n_{0}}}_{A}\underbrace{\sum_{i=1}^{n_{1}}{\frac{Y_{i}}{n_{1}}\hat{w}_{i}}}_{B}$
$=\frac{1}{n_{0}}\sum_{i=1}^{n_{1}}{Y_{i}\hat{w}_{i}}\rightarrow\mu_{0}^{1}.$ (7)
Note that $\textrm{E}(S_{i}w_{i})=n_{0}/n_{1}$ implies that $\textrm{E}\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}{w_{i}}\right)=n_{0}/n_{1}$, which implies that $\textrm{E}\left(\sum_{i=1}^{n_{1}}{w_{i}}\right)=n_{0}$. As such, the stabilized IOW estimator in Equation 4 is also consistent for the ATC.
3.3 Entropy balancing (matching-adjusted indirect comparison)
In Section 3.2, we described a “modeling approach” to weighting.66, 67 Propensity scores are explicitly modeled as a function of baseline covariates by a logistic regression, and are estimated by maximizing the fit of such regression, with inverse odds weights derived from the estimated propensity scores. This approach has certain limitations:66, 67
  • The estimated weights do not produce adequate covariate balance if the propensity score model is misspecified, and even a correctly specified model does not guarantee balance in finite samples;
@@ 3.3 Entropy balancing (matching-adjusted indirect comparison)
Such limitations motivate alternative “balancing” or “calibration” approaches to weighting. These estimate weights under the condition that covariates are balanced, viewing balance as an optimization problem, without explicitly modeling the propensity score. Generally, balancing approaches to weighting are: (1) less susceptible to bias by directly enforcing covariate balance; (2) produce more stable weights, which translate into larger effective sample sizes and more precise treatment effect estimation; and (3) are applicable where only aggregate-level marginal covariate moments are available for the external control.66, 67, 68, 69
Our focus here is on an entropy balancing approach37 called matching-adjusted indirect comparison (MAIC),38, 39, 40 but see Chattopadhyay et al (2020) and Filla et al (2024) for details about similar balancing techniques.66, 67 MAIC is the most popular balancing method in the context of externally controlled SATs. It has many features that are considered desirable: “linear double robustness”, minimally dispersed weights, and the estimation of odds weights that are guaranteed to be positive, resulting in increased interpretability and sample-boundedness, i.e., interpolating the observed data as opposed to extrapolating beyond its support.67, 70 We review the main steps of MAIC, building on prior literature.38, 69, 71, 72, 73, 74
@@
$\ln(\omega_{i})\propto\ln\Bigg{(}\frac{(1-e_{i})}{e_{i}}\Bigg{)}=\gamma_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma},$ (8)
where $\omega_{i}$ is a weight proportional to the inverse conditional odds of SAT participation for subject $i=1,\dots,n$, $\gamma_{0}$ is an intercept term parameter and $\mathbf{\gamma}$ is a vector of model parameters. In Equation 8, it is the log-odds of external control participation, $\operatorname{logit}(1-e_{i})$, that are linear across the covariate balance functions. Because $\operatorname{logit}(1-e_{i})=-\operatorname{logit}(e_{i})$, this implies that the log-odds of SAT participation, $\operatorname{logit}(e_{i})$, also vary linearly with $\mathbf{c}(\mathbf{X}_{i})$ as per Equation 2.
Signorovitch et al38 propose using the “method of moments” to estimate the model in Equation 8, such that:
$\frac{\sum_{i=1}^{n_{1}}\omega_{i}\mathbf{c}(\mathbf{X}_{i})}{\sum_{i=1}^{n_{1}}\omega_{i}}$ $=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{c}(\mathbf{X}_{i})$ (9)
[…]
@@
$Q(\hat{\mathbf{\gamma}})=\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right),$ (13)
as the derivative of $Q(\hat{\mathbf{\gamma}})$ with respect to $\hat{\mathbf{\gamma}}$ is $\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)\mathbf{c^{*}}(\mathbf{X}_{i})$. The objective function in Equation 13 is strictly convex and can be minimized using standard Newton-type convex optimization algorithms,37 yielding an unique finite solution corresponding to the global minimum of $Q(\hat{\mathbf{\gamma}})$.
We have $\omega_{i}\propto\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)\propto\exp\left(\left(\mathbf{c}(\mathbf{X}_{i})-\mathbf{\theta}\right)^{\top}\mathbf{\gamma}\right)=\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)$. Subject to the normalization constraint $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}=1$, weights for each individual $i=1,\dots,n_{1}$ in the SAT are estimated as:
$\hat{\omega}_{i}=\frac{\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}.$ (14)
@@
Note that this definition of the entropy balancing weights coincides with the definitions presented by Jiang et al (2024) 74 and by Jackson et al (2021).73
@@
Similar to Equation 3 and Equation 4, the ATC is estimated by contrasting mean absolute outcomes on the additive scale:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (15)
@@
where the weights have already been normalized to sum to one. Alternatively, fitting a weighted univariable regression of outcome on treatment to the concatenated IPD (with weights $\hat{\omega}_{i}$ for $i=1,\dots,n_{1}$ and unit weights for $i=n_{1}+1,\dots,n$) has been proposed, with the treatment coefficient of the fitted model yielding an ATC estimate.47, 52, 75, 76, 77 We discourage this approach. The weights estimated by Equation 14 are relative; their arbitrary rescaling by a constant of proportionality, e.g., a normalization constant, will also balance the specified covariate functions and does not affect $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$ in Equation 15.71, 73 Nevertheless, rescaling the weights while retaining weights of one for the external control subjects results in different fitted models.73
@@
The method of moments MAIC estimator gives the Lagrangian dual solution to an entropy balancing primal problem: minimizing the negative entropy of the weights.40, 42, 71 Namely, minimizing the objective function in Equation 13 is equivalent to minimizing the negative entropy $\sum_{i=1}^{n_{1}}\omega_{i}\ln(\omega_{i})$ with Lagrange multipliers.42, 71 The dual optimization problem is easier to solve than the primal37 and has been formulated in different ways.40, 71, 74 These may perform differently computationally – e.g., if minimization is performed on the log scale71 – but result in the same unique dual solution, up to numerical error, due to strict convexity. Because the negative entropy measures the distance of the weights from a uniform distribution, its minimization should produce less disperse weights than the modeling approach in Section 3.2.
While the entropy balancing weights ($\hat{\omega}_{i}$) defined in Equation 14 will be different than the IOW weights ($\widehat{w}_{i}$) obtained from maximum-likelihood estimation of the logistic regression model in Equation 2, $\hat{\nu}_{i}=n_{0}\hat{\omega}_{i}$ will consistently estimate the true inverse odds if the logistic regression model is correctly specified. To be clear, if the logistic regression model correctly specifies the true propensity score model, then we have both: $\widehat{w}_{i}\rightarrow\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$ (i.e., the IOW weights are consistent) and $\widehat{\nu}_{i}\rightarrow\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$ (i.e., the normalized entropy balancing weights are consistent), for $i$ in $n_{1}$; see Zhao and Percival (2017) 43 for details. Therefore, following the same logic as detailed in Section 3.2 for the IOW estimators, the entropy balancing estimator, as defined in Equation 15, is also consistent if the implied propensity score model is correctly specified.

### v2

@@ 3.1 Data and assumptions
As per Section 2, let $T=t$ denote a time-fixed binary treatment, with $t\in\{0,1\}$, such that $T=1$ represents the active intervention and $T=0$ the control. Let $S=s$ denote the data source, with $s\in\{0,1\}$, such that $S=1$ represents the SAT and $S=0$ the external data source. In addition, let $\mathbf{X}$ denote vector-valued pre-treatment baseline covariates, e.g., clinical or demographic characteristics, measured across the SAT and the external data source. Let $Y$ denote the clinical outcome of interest. We assume that only distributional differences in $\mathbf{X}$ are preventing exchangeability between the SAT and external subjects, and that covariates and outcomes are defined and measured similarly across data sources.
The observed IPD consist of $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$, $i=1,\dots,n_{1}+n_{0}$, realizations of $(S,\mathbf{X},T,Y)$ denoting the data source, baseline covariates, treatment assignment and observed outcome for subject $i$. Here, the SAT and external data source have been stacked, with $n_{1}$ and $n_{0}$ as the sample sizes of the SAT and the external data source, respectively. It is assumed that all individuals in the SAT are under $T=1$ and all individuals in the external data source are under $T=0$, such that the control group is fully external. To be clear, we have $S_{i}=T_{i}$ for all $i=1,\dots,n$, where $n=n_{1}+n_{0}$; and also have that $S_{i}=1$ and $T_{i}=1$ for all $i=1,\dots,n_{1}$, and $S_{i}=0$ and $T_{i}=0$ for all $i=n_{1}+1,\dots,n$. We shall assume that there is no missingness or measurement error.
The observed outcome for subject $i$ is $Y_{i}=Y_{i}^{1}T_{i}+Y_{i}^{0}(1-T_{i})$, where $Y_{i}^{t}$ is the potential outcome had subject $i$ been assigned treatment $t\in\{0,1\}$, with $Y_{i}=Y_{i}^{1}$ if $i=1,\dots,n_{1}$ and $Y_{i}=Y_{i}^{0}$ if $i=n_{1}+1,\dots,n$. Namely, the observed outcome for an individual in the SAT equals their potential outcome under the active intervention, and the observed outcome for an individual in the external data source equals their potential outcome under the control. Implicit in the notation is the stable unit treatment value assumption (SUTVA): that there is no interference between subjects and there is treatment version irrelevance, i.e., one well-defined version of the active intervention and the control across all subjects and data sources.zhou2024causal Also implicit is that there is no direct effect of trial participation.zhou2024causal Namely, that trial participation – in the SAT or a historical trial, for that matter – does not affect the outcome except through treatment assignment itself, i.e., there are no Hawthorne effects.braunholtz2001randomized, dahabreh2019hawthorne
@@
Two causal identification conditions, together known as strong ignorability, are required to construct a valid estimator of $\mu_{0}^{1}$. These ensure that the SAT and external control outcomes are comparable given adjustment for baseline covariates. The first assumption is conditional data source ignorability; formally, $Y_{i}^{1}\perp\!\!\!\!\perp S_{i}\mid\mathbf{X}_{i}$ for all $i=1,\dots,n$. Namely, conditional on baseline covariates, the potential outcome under the active intervention is independent of the data source. This is akin to the conditional constancy or exchangeability of absolute outcomes invoked for unanchored ITCs, used to transport mean absolute outcomes under $T=1$ from $S=1$ to $S=0$.phillippo2016nice, phillippo2018methods Conditional ignorability is a strong assumption, resting on the SAT and the external control capturing all variables that are prognostic of outcome under the active intervention.
The second assumption is positivity or overlap. That is, the support of the baseline covariates in the external control is contained within that of the SAT. Mathematically, the probability of SAT participation, conditional on the covariates necessary to ensure ignorability, should be bounded away from zero and one: $0<\textrm{Pr}(S=1\mid\mathbf{X}=\mathbf{x})<1$ for all $\mathbf{x}$ with positive density in the external control, i.e., for all $\mathbf{x}$ such that $f(\mathbf{x}\mid S=0)>0$. Hence, it is possible to have SAT subjects in all regions of the covariate space in $S=0$.zhou2024causal, dahabreh2019generalizing, buchanan2018generalizing Positivity violations can be deterministic or random. The former arise structurally, due to non-overlapping SAT and external control eligibility criteria. The latter arise empirically due to chance, particularly with small sample sizes.westreich2010invited To enforce positivity, analysts may subset the SAT based on the selection criteria of the external control.phillippo2016nice, phillippo2018methods However, this further reduces the sample size of the SAT. Positivity is typically assessed by comparing the empirical distributions of the covariates in the SAT and the external control.glimm2022geometric While outcome modeling-based approaches such as G-computation can overcome failures of positivity, they do so by potentially problematic and difficult-to-diagnose model-based extrapolation. Even minor model misspecification over the observed covariate space in the SAT may lead to poor extrapolation in unobserved regions of the covariate space.vo2023cautionary
Analogously, targeting the ATT would require constructing estimators for $\mu^{1}_{1}=\textrm{E}(Y^{1}\mid S=1)$ and $\mu^{0}_{1}=\textrm{E}(Y^{0}\mid S=1)$. Here, the challenge is the estimation of $\mu^{0}_{1}$ because outcomes under the control have not been generated in the SAT. The conditional ignorability assumption would formally be $Y_{i}^{0}\perp\!\!\!\!\perp S_{i}\mid\mathbf{X}_{i}$ for all $i=1,\dots,n$, and would rest on the SAT and the external control measuring all variables that are prognostic of outcome under the control. The positivity assumption would be $0<\textrm{Pr}(S=0\mid\mathbf{X}=\mathbf{x})<1$ for all $\mathbf{x}$ with positive density in the SAT, $f(\mathbf{x}\mid S=1)>0$, such that the support of the baseline covariates in the SAT is contained within that of the external control and it is possible to have external control subjects in all regions of the SAT covariate distribution. In this setting, analysts may apply the SAT selection criteria to the external control to guarantee that there is sufficient overlap. While the methods in the next sections target the ATC, we present any required modifications to target the ATT in the Supplementary Material.
@@ 3.2 Inverse odds weighting
We first present a covariate adjustment method that models the data source assignment mechanism, conditional on baseline covariates, to estimate weights.ren2023comparing, loiseau2022external, josey2021transporting, dahabreh2020extending, colnet2024causal Where the target estimand is the ATC, SAT subjects are weighted by their inverse conditional odds of SAT participation – their conditional odds of external control participation – to transport the SAT outcomes to the external control (sub) population. Such “inverse odds” weights (IOW) are defined as:
@@
where $\operatorname{logit}(e_{i})=\ln\left((e_{i})/(1-e_{i})\right)$, $\alpha_{0}=\ln\left(\textrm{Pr}(S_{i}=1|\mathbf{c}(\mathbf{X}_{i})=0)/(\textrm{Pr}(S_{i}=0|\mathbf{c}(\mathbf{X}_{i})=0)\right)$ is an intercept term, $\mathbf{\alpha}$ is a vector of regression parameters, and $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i}),\dots,c_{k}(\mathbf{X}_{i})]^{\top}$ is a vector of covariate “balance functions” for subject $i=1,\dots n$. This is the set of functions containing the distributional features to be balanced between the SAT and the external control,josey2021transporting, josey2022calibration potentially including sensible transformations of the covariates, e.g., polynomials and interaction terms.
[…]
The ATC is estimated by contrasting the weighted average of observed outcomes under the active intervention with the unweighted average of observed outcomes for the external control. As per Section 2, mean absolute outcomes are converted to the additive scale imposed by link function $g(\cdot)$ prior to taking the difference between treatments on such scale, leading to the IOW estimator for the ATC:
@@
The mean absolute outcome estimate for the active intervention can be bounded within its feasible range, e.g., between 0 and 1 for probabilities, by normalizing or “stabilizing” the weights so that they sum to one.dahabreh2019generalizing, dahabreh2020extending This results in the alternative normalized IOW estimator for the ATC:
@@
which should provide improved finite sample properties and more stable and precise estimation.busso2014new, hernan2000marginal Drawing an analogy with survey sampling, the estimator in Equation 3 is a Horvitz-Thompson-type estimator and that in Equation 4 is a Hajek-type estimator, with the latter typically considered to improve the performance of the former in the literature.sarndal2003model
[…]
Unfortunately, this “modeling approach” to weighting,chattopadhyay2020balancing, filla2024balancing where propensity scores are explicitly modeled as a function of baseline covariates by a logistic regression, then estimated by maximizing the fit of such regression, has certain limitations:chattopadhyay2020balancing, filla2024balancing
  • The resulting weights do not produce adequate covariate balance if the propensity score model is misspecified, and even a correctly specified model does not guarantee balance in finite samples;
@@
3.3 Entropy balancing (matching-adjusted indirect comparison)
The limitations of the weighting methods in Section 3.2 motivate alternative “balancing” or “calibration” approaches to weighting. These estimate weights under the condition that covariates are balanced, viewing balance as an optimization problem, without explicitly modeling the propensity score. Generally, balancing approaches to weighting are: (1) less susceptible to bias by directly enforcing covariate balance; (2) produce more stable weights, which translate into larger effective sample sizes and more precise treatment effect estimation; and (3) are applicable where only aggregate-level marginal covariate moments are available for the external control.chattopadhyay2020balancing, filla2024balancing, hirshberg2017two, wang2021matching
Our focus here is on an entropy balancing approachhainmueller2012entropy called matching-adjusted indirect comparison (MAIC),signorovitch2010comparative, cheng2020statistical, josey2021transporting but see Chattopadhyay et al (2020) and Filla et al (2024) for details about similar balancing techniques.chattopadhyay2020balancing, filla2024balancing MAIC is the most popular balancing method in the context of externally controlled SATs. It has many features that are considered desirable: “linear double robustness”, minimally dispersed weights, and the estimation of odds weights that are guaranteed to be positive, resulting in increased interpretability and sample-boundedness, i.e., interpolating the observed data as opposed to extrapolating beyond its support.filla2024balancing, tan2010bounded We review the main steps of MAIC, building on prior literature.signorovitch2010comparative, wang2021matching, phillippo2020equivalence, remiro2022two, jackson2021alternative, jiang2024comprehensive
@@ 3.3 Entropy balancing (matching-adjusted indirect comparison)
$\ln(v_{i})\propto\ln\Bigg(\frac{(1-e_{i})}{e_{i}}\Bigg)=\gamma_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma},$ (5)
where $v_{i}$ is a weight proportional to the inverse conditional odds of SAT participation for subject $i=1,\dots,n$, $\gamma_{0}$ is an intercept term parameter and $\mathbf{\gamma}$ is a vector of model parameters. In Equation 5, it is the log-odds of external control participation, $\operatorname{logit}(1-e_{i})$, that are linear across the covariate balance functions. Because $\operatorname{logit}(1-e_{i})=-\operatorname{logit}(e_{i})$, this implies that the log-odds of SAT participation, $\operatorname{logit}(e_{i})$, also vary linearly with $\mathbf{c}(\mathbf{X}_{i})$ as per Equation 2.
Signorovitch et alsignorovitch2010comparative propose using the “method of moments” to estimate the model in Equation 5, such that:
$\frac{\sum_{i=1}^{n_{1}}v_{i}\mathbf{c}(\mathbf{X}_{i})}{\sum_{i=1}^{n_{1}}v_{i}}$ $=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{c}(\mathbf{X}_{i})$ (6)
$\frac{\cancel{\exp\left(\gamma_{0}\right)}\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)\mathbf{c}(\mathbf{X}_{i})}{\cancel{\exp\left(\gamma_{0}\right)}\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)}$ $=\underbrace{\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{c}(\mathbf{X}_{i})}_{\hat{\mathbf{\theta}}}$ (7)
@@
[…]
$\hat{v}_{i}=\frac{\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}$. (11)
@@
Note that this definition of the entropy balancing weights coincides with the definitions presented by Jiang et al (2024) jiang2024comprehensive and by Jackson et al (2021).jackson2021alternative
@@
Similar to Equation 4, the MAIC estimator for the ATC contrasts absolute outcomes on the additive scale:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{v}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (12)
@@
where the weights have already been normalized to sum to one. Alternatively, fitting a weighted univariable regression of outcome on treatment to the concatenated IPD (with weights $\hat{v}_{i}$ for $i=1,\dots,n_{1}$ and unit weights for $i=n_{1}+1,\dots,n$) has been proposed, with the treatment coefficient of the fitted model yielding an ATC estimate.ishak2015simulation, park2024unanchored, jiang2020performance, hatswell2020effects, chandler2024uncertain We discourage this approach. The weights estimated by Equation 11 are relative; their arbitrary rescaling by a constant of proportionality, e.g., a normalization constant, will also balance the specified covariate functions and does not affect $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$ in Equation 12.phillippo2020equivalence, jackson2021alternative Nevertheless, rescaling the weights while retaining weights of one for the external control subjects results in different fitted models.jackson2021alternative
@@
The method of moments MAIC estimator gives the Lagrangian dual solution to an entropy balancing primal problem: minimizing the negative entropy of the weights.josey2021transporting, cheng2023double, phillippo2020equivalence Namely, minimizing the objective function in Equation 10 is equivalent to minimizing the negative entropy $\sum_{i=1}^{n_{1}}v_{i}\ln(v_{i})$ with Lagrange multipliers.cheng2023double, phillippo2020equivalence The dual optimization problem is easier to solve than the primalhainmueller2012entropy and has been formulated in different ways.josey2021transporting, phillippo2020equivalence, jiang2024comprehensive These may perform differently computationally – e.g., if minimization is performed on the log scalephillippo2020equivalence – but result in the same unique dual solution, up to numerical error, due to strict convexity. Because the negative entropy measures the distance of the weights from a uniform distribution, its minimization should produce less disperse weights than the modeling approach in Section 3.2.